A local pizza shop operates 7 days a week due to its popularity amongst customers. The pizza shop has an annual demand of 1500 cans of tomato sauce. The cost per can is $10, the annual holding cost of inventory is 12% of cost and the administrative costs to place an order is $50.

QUESTION:

A) What is the optimal order quantity (round to the next whole value) would be to minimize the pizza shop’s annual inventory cost?
Answer #1

DEMAND = 1500
ORDERING COST = 50
HOLDING COST = 1.2
COST PER UNIT = 10

EOQ = SQRT(2 * D * S / H), WHERE D = DEMAND, S = ORDERING COST AND H = HOLDING COST = SQRT(2 * 1500 * 50 / 1.2) = 354

Do leave a like for the answer if you found it helpful.